OpenStudy (anonymous):

lim (sin x )^tanx as x->0 ?

OpenStudy (agent0smith):

I think you can apply l'hopitals rule, after modifying it a bit...

OpenStudy (anonymous):

Yeah,just solved it with these rules,it's equal 1.

OpenStudy (agent0smith):

Yep :)

OpenStudy (agent0smith):

I had to try it myself... I'm guessing you ended up with something like \[\Large \lim_{x \rightarrow 0} e^{-\cos x \sin x}\]which if you plug in 0 gives you 1.

OpenStudy (agent0smith):

After modifying it using logs, then using l'hopitals rule.

OpenStudy (anonymous):

Can you please write the whole way of it ? I did it a bit different ...

OpenStudy (agent0smith):

lim (sin x )^tanx as x->0 First let y = sinx^tanx, and take logs of both sides \[\large \ln y = \tan x * \ln (\sin x)\] put both sides to the power of e \[\LARGE y = e^{\tan x *\ln (\sin x)}\]

OpenStudy (agent0smith):

So that was just to modify the limit. Then write it as \[\LARGE \lim_{x \rightarrow 0 } e^{\tan x *\ln (\sin x)}\] for which you can bring the limit up into the exponent \[\huge e^{ \left( \lim_{ x \rightarrow 0}\tan x *\ln (\sin x) \right)} \]

OpenStudy (agent0smith):

Site crashed... that's why i write solutions in pieces :P To apply l'hopitals rule, get the tanx into the denominator so that it's in inf/inf form: \[\huge e^{ \left( \lim_{ x \rightarrow 0}\frac{ \ln (\sin x) }{ \tan^{-1} x } \right)}\] cos now if you plug in 0 you get -inf/inf

OpenStudy (agent0smith):

Now you can differentiate using hopitals \[\huge e^{ \left( \lim_{ x \rightarrow 0}\frac{ \frac{ \cos x }{\sin x }}{ - \cot^2 x } \right)}\] which after simplifying gives what i gave above.

OpenStudy (anonymous):

Thank you very much ! :)

OpenStudy (agent0smith):

No prob, i'm just satisfied i could actually solve it! hehe

OpenStudy (zarkon):

\[\tan^{-1}(0)=0\]

OpenStudy (zarkon):

oh..i see what you did...you should not use that notation

OpenStudy (agent0smith):

ohh, yes. My bad.

OpenStudy (agent0smith):

But it's not too hard to figure out what i meant in this case, since i mentioned putting it into the denominator.

OpenStudy (zarkon):

if you want 1 over the tangent you should write \[\frac{1}{\tan(x)}\] or \[\left[\tan(x)\right]^{-1}\] \[tan^{-1}(x)\] typically means the inverse tangent

OpenStudy (agent0smith):

haha, i know :P

OpenStudy (agent0smith):

I wrote it as (tanx)^-1 and then thought "no, tan^-1 x is neater..."

OpenStudy (zarkon):

I would have known that if i actually read from the beginning and not just your last post ;)

OpenStudy (agent0smith):

haha :)

OpenStudy (zarkon):

cheers!